The sum of two consecutive odd integers is -40. Find the integers. please help meeee
ASAP this is a homework.!

Answer by stanbon(75887)

sum of two consecutive odd integers is -40.
----------------
1st: 2x-1
2nd: 2x+1
----
sum = -40
4x = -40
---
x = -10
----
1st: 2x-1 = -20-1 = -21
2nd: 2x+1 = -20+1 = -19
============================
Cheers,
Stan H.